# mCLM: A Function-Infused and Synthesis-Friendly Modular Chemical Language Model

### v1

@@ 2.4 Chemical-Language Modeling
Figure 2 illustrates the architecture of mCLM. The model is a sequential generative model that processes molecule and text sequences in a unified manner. It adopts a Transformer architecture, which is well-suited for handling sequential data and allows for using pre-trained language models as a backbone. After tokenizing the molecules as mentioned in the last section, we encode each building block using graph neural networks (GNNs) [19, 83, 26]. These representations are then concatenated with natural language embeddings at positions where molecule entity names appear. This results in a form of “code-switched” language which blends molecular structures with natural language descriptions. The feature sequence is then fed into a Transformer decoder-only architecture, which predicts the next token based on previous tokens. This allows for pre-training the model on a large corpus of multi-modal data, enabling it to learn and “talk about” the relationships between the different modalities and their respective representations.
We train this model—mCLM—on top of open-source large language models pretrained on general-domain corpora, allowing us to leverage their linguistic capabilities without incurring the computational cost of training from scratch. As the training objective, we adopt a unified categorical cross-entropy (CCE) loss applied to both natural language and molecular tokens:
$\mathcal{L}=H(P(\mathbf{X}),P_{\theta}(\mathbf{X})),~{}~{}~{}\text{logit}(P_{\theta}(v\mid\mathbf{X}_{1\cdots i-1}))=\begin{cases}\mathbf{c}^{\top}\mathbf{e}_{v},v\in\mathcal{V}_{\text{natural language}}\\ \mathbf{c}^{\top}f_{\theta}(\text{GNN}_{\theta}(v)),v\in\mathcal{V}_{\text{molecular building block}}\\ \end{cases}$
Specifically, the loss is computed between the ground truth distribution $P(\mathbf{X})$ and the model-predicted distribution $P_{\theta}(\mathbf{X})$ over the combined vocabulary of natural language and molecular building blocks. The model generates logits for the next token $v$ by computing the dot product between the contextual representation $\mathbf{c}$ with token embeddings. $\mathbf{c}$ is produced by the Transformer given the previous tokens $\mathbf{X}_{1\cdots i-1}=[X_{1},\cdots,X_{i-1}])$. For natural language tokens, the embedding $\mathbf{e}_{v}$ is directly taken from the pretrained natural language model. The embedding for molecular building blocks is computed by passing the building block’s graph through a GNN, followed by a linear adapter function $f_{\theta}$ to project it into the same embedding space. This formulation enables joint training over both modalities using a single loss function. More detailed training procedure is described in Appendix C.
@@ 2.5 Critical Chemical Reasoning
Chemical reasoning over molecules often involves optimization of multiple functions, such as toxicity, bioactivity, binding affinity, etc. Therefore, it is not a straightforward task to propose an ideal molecule structure, especially with only a single attempt. In nature, organisms evolved over billions of years to produce complex molecules that are optimal for their use. Additionally, many functions positively or negatively correlate with each other. Optimizing one function may lead to trade-offs in others. For example, many drugs with higher potency were rejected due to increased toxicity to patients. To address this, we propose a reasoning process that allows the model to refine its own generated molecules and iteratively improve their desired functions. At each step, mCLM proposes a modification of the molecule for improvement by considering one or more functions and self-evaluates the effect of the proposed modification. Then in the next iteration, mCLM identifies those functional building blocks that still have room for improvement. This process is repeated until a maximum number of iterations is reached, or if mCLM fails to find additional reasonable modifications. As illustrated in Figure 2, this process can be summarized in the following algorithm:
@@
8:     end if
@@
9:     mCLM generates a candidate modification $M^{\prime}$ by replacing, adding, or removing building blocks in $M$
10:     mCLM evaluates $M^{\prime}$ with respect to desired functions
11:end while
3 Experimental Evaluation
3.1 Creating Oracle Models for Evaluation
To evaluate the performance of our generated molecules, we construct oracle models focused on Absorption, Distribution, Metabolism, Excretion, Toxicity (ADMET) property prediction. We select 6 tasks from the Therapeutics Data Commons (TDC) benchmark [33]: AMES (mutagenicity), BBBP (blood-brain barrier permeability), CYP3A4 inhibition (metabolism), DILI (drug-induced liver injury), HIA (human intestinal absorption), and PGP (P-glycoprotein substrate classification).
Models are trained using predefined scaffold-based data splits with an 8:1:1 ratio for train, validation, and test sets. The detailed training procedure is described in Appendix B.2.3. We opt to use robust foundation models—FARM [71], ChemBERTA-2 [2], and a GNN [19]—for ensemble learning. To build the ensemble, we use each model as a feature extractor. The extracted features are concatenated and passed through a fully connected layer for final prediction.
3.2 Improving FDA-Approved Drugs with Out-of-Vocabulary Blocks
As a large-scale test, we apply the mCLM to improve all FDA-approved drugs consisting of at least 3 blocks. This amounts to 430 molecular structures and 796 unseen blocks. Since most of these molecules (426/430) contain blocks that were not present in the 1,000 blocks used for training, this presents an opportunity to find how the mCLM performs out-of-distribution. Results in Table 1 show that improvement is achieved for 5/6 properties, even though the model has not seen almost half of the blocks in its vocabulary during training.
Property FDA Drug Modified AMES Mut. ()($\downarrow$)( ↓ ) 59.5 54.0 CYP3A4 Inhib. ()($\downarrow$)( ↓ ) 2.0 1.2 BBBP ()($\uparrow$)( ↑ ) 37.6 41.4 HIA ()($\uparrow$)( ↑ ) 93.2 97.6 DILI ()($\downarrow$)( ↓ ) 66.2 55.5 PGP ()($\downarrow$)( ↓ ) 66.0 68.0
Table 1: Average pharmacokinetic and toxicity properties of FDA drugs with 3 or more blocks and their proposed modifications. ($\downarrow$: lower is better, $\uparrow$: higher is better; p < 0.05 by Wilcoxon Test [105, 74].)
3.3 Case Study: Multi-step Critical Reasoning to Resurrect the “Fallen Angels”
Figure 4: Examples of fallen angel property modification.
In a similar vein, there are many new drug candidates that almost make it to FDA approval but fall short for various reasons when being tested in clinical trials. For example, Evobrutinib is a Bruton’s tyrosine kinsase (BTK) inhibitor that went through clinical trial as a drug for relapsing Multiple Sclerosis. However, the FDA placed a partial clinical hold on Phase III trials in April 2023 after two patients showed signs of drug-induced liver injury [67]. TNG348 is a USP1 inhibitor designed for treating BRCA1/2-mutant and HRD cancers, but it failed in phase 1/2 clinical trials due to liver abnormalities [35, 80]. These “fallen angels” represent tremendous opportunities for impactful engagement of the AI/chemistry interface, because much is known about the strengths of each of these small molecules, and it is also known why they fell short. Fixing such fallen angels is a high leverage opportunity for the function-infused and synthesis-friendly mCLM to contribute. Herein we demonstrate this potential with some real-world examples of fallen angels that can be optimized with mCLM. Importantly, and uniquely, all of the new compounds proposed by mCLM are, by design, well-suited for rapid automated assembly using robots.
Figure 4 shows an application of the mCLM to these two molecules. For both, the inital step is to optimize DILI, the reason the drugs failed in clinical trials. Following that, the mCLM is applied to fix other properties which were made worse in the previous attempt (PGP for Evobrutinib and BBBP for TNG248). For good measure, another property of each molecule is then improved. Notably, at each step, the mCLM only makes minor modifications to each drug of roughly 1 building block. While the mCLM shows promising results for repairing these drugs, it is worth noting that drug discovery is a many-objective optimization problem. While we are able to generate proposed molecules with improved toxicity relative to Evobrutinib and TNG348, as well as other properties, yet other important properties may still have been compromised. Future work may want to investigate longer reasoning chains across a wider variety of properties.
3.4 Library Design: Identifying Functionally-Important Blocks
One notable thing about the architecture of the mCLM is the ability to identify blocks which are preferred for certain functions. This can help inform virtual screening campaign design (e.g., which 10 blocks should we use to get the best chemical space for BBBP?), and it can also be useful for stimulating scientific inquiry. As an example, Figure 5 shows the most frequent modifications preferred by the mCLM for improving DILI in FDA drugs. Interestingly, these modification tend to be small changes from the original molecule block.
4 Related Work
4.1 Applying LLMs to Chemical Representation and Discovery
Following their success in general natural language processing tasks [123], large language models (LLMs) have been increasingly adopted for solving a wide range of problems in computational chemistry [116, 119]. Existing efforts can be broadly categorized into two groups based on molecular representation: (1) SMILES-based modeling [24, 61, 72, 124], and (2) graph-based modeling [59]. For SMILES-based approaches, MolT5[18] introduces a multimodal language model that jointly understands and generates textual and SMILES representations. GIMLET[122] unifies graph and textual information using a distance-aware graph-text Transformer. LlaSMol[115] is instruction-tuned for property prediction and molecular description. DrugAssist [114] explores interactive molecule optimization by enabling LLMs to perform goal-driven modifications through human-in-the-loop dialogue. Graph-based approaches aim to capture molecular structure more explicitly. MolCA [59] integrates graph encoders with Galactica [89] using a cross-modal projector. 3D-MoLM[53] extends this paradigm to 3D molecular structures by incorporating a 3D molecular encoder. Llamole[55] further advances this direction by enabling interleaved generation of text and molecular graphs, supporting applications such as retrosynthetic planning and molecular inverse design. However, most molecule language models tokenize at the level of atoms [99, 2, 126, 109] and/or they do not infuse synthesis considerations up front, which precludes matching tokens with desired functions and/or leads to AI- generated new molecular targets that are not readily synthesizable.
4.2 Functional Groups/Fragments
Functional groups/fragments are core structural motifs that govern molecular reactivity, biological activity, and physicochemical properties, forming the basis of chemical understanding [23, 75, 39]. Their study has propelled advances in medicinal chemistry, natural product research, and material science by guiding synthesis and performance optimization [23, 22, 3, 5].
Given their interpretability and functional relevance, functional groups/fragments are also increasingly being incorporated into machine learning models for molecular representation [46, 70, 29, 38, 68, 102, 118, 120]. Some approaches employ a masking strategy for masked atom prediction by simultaneously masking all atoms within a functional group, applied to both molecular graph and SMILES representations [103, 46]. Other approaches include augmenting SMILES strings with functional group annotations [68], building functional group-level graphs [68], or leveraging hierarchical architectures that model both atomic-level and functional module-level representations [29, 70].
Our work goes beyond mere chemical representation by attempting to encode molecules at the building-block level within natural language LLMs, paving the way for chemical reasoning. These methods also often overlook the synthesizability of molecules. They tend to break molecules at bonds that are difficult to form through chemical reactions, resulting in generated molecules that are challenging—or even impossible—to synthesize using standard reaction pathways. This gap limits the practical applicability of the learned representations, especially for tasks such as molecular generation or optimization.
4.3 Multimodality, Multilinguality and Code-Switching in LMs
This work is also inspired by exploration in enabling LLMs to understand multiple languages [34, 10, 117, 15, 91] or data modalities such as vision and audio [52, 1, 57, 64, 98, 96, 113, 13, 101, 12, 88, 111]. [32] fed genome sequences to LLMs and found them capable of suggesting common functions for gene sets while also providing interpretable rationales. Explorations on proteins [60, 78] show success in protein-to-text and text-to-protein generation in applications including antibody design and protein understanding. This work is also related to other recent studies that have explored the integration of graph neural networks (GNNs) as feature vectors into LMs capture both textual and structural information [59, 122, 37].
5 Conclusions and Future Work
In this work, we propose the modular Chemical-Language Model, which is the first attempt to jointly model natural language sequences with modular chemical language. mCLM is capable of conducting critical chemical reasoning by iteratively refining molecule designs over diverse objectives and chemical knowledge. Future work could incorporate richer information such as 3D structures of molecules and physical constraints into mCLM representation. Other modalities such as protein and nucleic acid sequences will further enable mCLM to reason on biological activity, protein docking knowledge, and individuals’ genetic profiles. Future efforts could also extend chemical reasoning to additional aspects such as collecting and leveraging results from physical simulation tools, including protein interaction dynamics, retrieval from chemical and reaction knowledge bases, etc. Another important direction is to introduce multi-agent interactions with other AI agents, human scientists, and automatic laboratories, where proposed molecules can be efficiently synthesized using modular chemistry technologies. There is also great potential in including alternative scientific hypothesis generation and plausibility prediction. In the long-term, we envision a landscape where mCLMs can be included in a more comprehensive chemical research workflow consisting of fully automated and never-ending reasoning - proposal - synthesis - physical testing - feedback - reasoning loops. mCLM can serve as a preliminary step in bridging AI and physical world experiments, shattering the barriers that currently preclude non-specialists in organic synthesis from meaningfully participating in small molecule innovation.
Acknowledgments and Disclosure of Funding
We would like to thank Kyunghyun Cho, Anna Hart, Hyeonjeong Ha, Gabriele Scalia, Yanru Qu, and Jeonghwan Kim for helpful discussion. This research is based upon work supported by the Molecule Maker Lab Institute: an AI research institute program supported by NSF under award No. 2019897, and IBM-Illinois Discovery Accelerator Institute (IIDAI) Center. The core idea of decomposing-and-reassembling functional parts into novel concepts is inspired by research supported by the U.S. DARPA ECOLE Program No. HR00112390060. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect those of the National Science Foundation.
Appendix A Pretraining Data Mixture

### v2

@@ 2.2 Chemical-Language Modeling
Figure 2 illustrates the architecture of mCLM. The model is a sequential generative model that processes molecule and text sequences in a unified manner. It adopts a Transformer architecture, which is well-suited for handling sequential data and allows for using pre-trained language models as a backbone. After tokenizing the molecules as mentioned in the last section, we encode each building block using graph neural networks (GNNs; Edwards et al., 2024a; Sprueill et al., 2024; Gasteiger et al., 2021) before passing them through an adaptor module. These representations are then concatenated with natural language embeddings at positions where molecule entity names appear. This results in a form of “code-switched” language which blends molecular structures with natural language descriptions. The feature sequence is then fed into a Transformer decoder-only architecture, which predicts the next token based on previous tokens. This allows for pre-training the model on a large corpus of multi-modal data, enabling it to learn and “talk about” the relationships between the different modalities and their respective representations.
We train mCLM on top of open-source pretrained large language models (we adopt Qwen2.5 Yang et al. (2024) as the starting point), allowing us to leverage their natural language understanding and reasoning capabilities without incurring the computational cost of training from scratch. As the training objective, we adopt a unified categorical cross-entropy (CCE) loss applied to both natural language and molecular tokens:
[…]
Specifically, the cross-entropy loss is computed between the ground truth distribution $P(\mathbf{x})$ and the model distribution $P_{\theta}(\mathbf{x})$ over the combined vocabulary of natural language and molecular building blocks. The model generates logits for the next token $v$ by computing the dot product between the contextual representation $\mathbf{c}_{i}$ with token embeddings. $\mathbf{c}_{i}$ is produced by the Transformer given the previous tokens $\mathbf{x}_{1\cdots i-1}=[X_{1},\cdots,X_{i-1}])$. For natural language tokens, the embedding $\mathbf{e}_{v}$ is directly taken from the pretrained natural language model. The embedding for molecular building blocks is computed by passing the building block’s graph through a GNN, followed by a linear adapter function $f_{\psi}$ to project it into the same embedding space. This formulation enables joint training over both modalities using a single loss function. We train on approximately 1 million examples consisting of the 1,000 most frequent molecular building blocks, sampled from a new dataset we constructed. This dataset contains over 36 million molecular instructions, 1,000 tasks, and 8 million molecules. More details on the training procedure are described in Appendix E. Notably, the architecture of mCLM is compatible with larger vocabularies, and we anticipate future scaling studies as computational resources permit.
@@ 2.3 Critical Chemical Reasoning
Chemical reasoning over molecules often involves optimizing multiple functions, such as toxicity, bioactivity, and binding affinity. Therefore, it is not a straightforward task to propose an ideal molecule structure, especially with only a single attempt. Optimizing one function may lead to trade-offs in others. For example, many drugs with higher potency were rejected due to increased toxicity to patients. To address this, we propose a reasoning process that allows the model to refine its own generated molecules and iteratively improve their desired functions. At each iteration, we evaluate the properties and identify one that still requires improvement, and mCLM proposes a modification of the molecule targeting this property. This process is repeated until a maximum number of iterations is reached. This process is summarized in Appendix Algorithm 1.
3 Experimental Evaluation
3.1 Creating Oracle Models for Evaluation
To evaluate the performance of our generated molecules, we construct oracle models focused on Absorption, Distribution, Metabolism, Excretion, Toxicity (ADMET) property prediction. We select 6 tasks from the Therapeutics Data Commons (TDC) benchmark (Huang et al., 2021): AMES (mutagenicity), BBBP (blood-brain barrier permeability), CYP3A4 inhibition (metabolism), DILI (drug-induced liver injury), HIA (human intestinal absorption), and PGP (P-glycoprotein substrate classification). The detailed training procedure is described in Appendix D.2.3.
3.2 Improving FDA-Approved Drugs with Out-of-Vocabulary Blocks
In practice, real-world drug discovery is often driven by the optimization of known molecules, as existing approved drugs offer a more direct and promising path to safe and effective new drugs. Therefore, we evaluate mCLM’s drug proposal capability in improving FDA-approved drugs. Specifically, we apply the mCLM to improve the 6 properties for all FDA-approved drugs consisting of synthesis-guaranteed blocks. This amounts to 122 molecular structures and 153 unseen blocks. We confine the output vocabulary of mCLM to 582 synthesis-guaranteed blocks. Most of these drugs (120/122) contain blocks that were not present in the 1,000-block training vocabulary, presenting an opportunity to examine how the mCLM works on out-of-distribution molecules. Results in Table 3.2 show that improvements are achieved for all properties [^2] We also test a different distribution of 430 non-synthesis-guaranteed FDA drugs in Appendix A.2.. This shows that mCLM is not limited to a fixed vocabulary but can generalize beyond its training coverage. We also compare our approach against a wide range of strong text-based molecule editing baselines, including MoleculeSTM (Liu et al., 2022), FineMolTex (Li et al., 2025b), GPT-4o, GPT-5, Gemini-2.5-Flash (Gemini-2.5-F), LDMol (Chang and Ye, 2024), and Claude 3.5 Haiku (Claude-3.5-H). Despite the fact that all of the baseline models are allowed to generate without synthesis guarantees, their average improvements fall behind mCLM.
Table 1: Average pharmacokinetic and toxicity properties of FDA drugs composed of synthesis-guaranteed blocks, as well as their proposed modifications. ($\downarrow$: lower is better, $\uparrow$: higher is better). Green = better than FDA, Red = worse, Light green bold = best overall per column.
Model AMES $(\downarrow)$ BBBP $(\uparrow)$ CYP3A4 $(\downarrow)$ DILI $(\downarrow)$ HIA $(\uparrow)$ PGP $(\downarrow)$ Avg. Improv. FDA Drug 47.8 61.4 2.1 60.1 98.96 64.6 0.00 % MoleculeSTM 47.1 63.4 2.2 59.3 98.76 64.1 0.31 % FineMolTex 47.5 66.0 2.4 59.5 98.84 64.0 -0.73 % LDMol 49.0 63.5 2.5 57.5 99.0 67.6 -3.07 % GPT-4o 46.0 72.1 2.2 60.8 99.3 65.2 2.45 % GPT-5 49.1 70.2 2.4 61.2 99.2 65.0 -0.82 % Claude-3.5-Haiku 49.3 72.2 2.2 58.6 98.3 65.7 1.64 % Gemini-2.5-Flash 45.0 72.5 1.8 58.1 99.1 64.5 6.98 % mCLM (Ours) 44.4 85.2 1.4 53.7 98.99 64.4 15.0 %
The key to expediting the drug creation process is to discover potent molecular candidates that are simultaneously synthesis-friendly. While mCLM shows strong property editing results, its key benefit lies in its synthesis-friendly nature. We assessed the synthesizability of generated molecules by computing synthetic accessibility (SA scores) (Ertl and Schuffenhauer, 2009) as a quick heuristic (see Table 3). Then, as a more rigorous assessment, we consider Allchemy , which is the state-of-the-art retrosynthesis software (Wołos et al., 2022; Strieth-Kalthoff et al., 2024c). Allchemy is computationally expensive, but it evaluates synthesizability to the best of publicly available human chemical knowledge. For example, it finds synthetic routes for 98.1% of the FDA-approved molecules. We select the top 3 models by SA score (mCLM, MolSTM, and FineMolTex) and randomly sample 200 generated molecules from each to be assessed by Allchemy on a supercomputing cluster.
Table 2: Synthetic accessibility (SA) (Ertl et al., 2009), validity, and retrosynthetic results across baselines. Synthesizability indicates a retrosynthetic route was found. Makeability =Valid $\times$ Synth.
Model SA Validity (%) Synthesizability (%) Makeability (%)
FDA 2.70 100.0 98.11 98.11
MoleculeSTM 2.64 93.80 91.03 85.39
FineMolTex 2.58 94.20 90.15 84.96
mCLM (Ours) 2.43 100.0 98.23 98.23
As shown in Table 2, molecules proposed by mCLM are 100% valid (syntactically correct) and 98.2% synthesizable, which is superior even to the FDA drugs. In contrast, MoleculeSTM outputs are valid only 93.9% of the time, and among those, only 90.3% are predicted to be synthesizable using exhaustive retrosynthesis search. Out of MoleculeSTM-generated molecules, only 85.4% can be made ($93.8\times 0.91=85.4$).
3.3 Library Design: Identifying Functionally-Important Blocks
As a by-product, mCLM is also able to identify blocks which are preferred for certain functions. This can help inform virtual screening campaign design (e.g., which 10 blocks should we use to get the best chemical space for BBBP?), and it can also be useful for stimulating scientific inquiry.
As an example, we calculated the most frequent modifications preferred by the mCLM to improve DILI in FDA drugs (Figure 4). Notably, modification 3 resembles a newly discovered modification for reducing amphotericin B toxicity by human scientists in Maji et al. (2023).
3.4 Case Study: Multi-step Reasoning to Resurrect the “Fallen Angels”
There are many new drug candidates that almost reach FDA approval but fall short for various reasons when tested in clinical trials. For example, Evobrutinib is a Bruton’s tyrosine kinsase (BTK) inhibitor that went through clinical trial as a drug for relapsing Multiple Sclerosis. However, the FDA placed a partial clinical hold on Phase III trials in April 2023 after two patients showed signs of drug-induced liver injury (Montalban et al., 2024). TNG348 is a USP1 inhibitor designed for treating BRCA1/2-mutant and HRD cancers, but it failed in phase 1/2 clinical trials due to liver abnormalities (Inc, 2024; Simoneau et al., 2025). These “fallen angels” represent tremendous opportunities for impactful engagement of the AI/chemistry interface, because much is known about the strengths of each of these small molecules, and it is also known why they fell short. Fixing such fallen angels is a high leverage opportunity for the function-infused mCLM to contribute.
Figure 5: Examples of fallen angel property modification.
Figure 5 shows an application of the mCLM, without synthesis restrictions on the vocabulary, to these two molecules. For both, the initial step is to optimize DILI, the reason the drugs failed in clinical trials. Following that, the mCLM fixes other properties which were made worse in the previous attempt (PGP for Evobrutinib and BBBP for TNG248). For good measure, another property of each molecule is then improved. Notably, at each step, the mCLM only makes minor modifications to each drug of roughly 1 building block. While the mCLM shows promising results for repairing these drugs, it is worth noting that drug discovery is a many-objective optimization problem. While we are able to generate molecules with improved toxicity relative to Evobrutinib and TNG348, as well as other properties, yet other important properties may still have been compromised. Future work may want to investigate longer reasoning chains across a wider variety of properties. For comparison, we show a (less-successful) version of this experiment using MoleculeSTM A.3.
4 Related Work
Protein language models (Lin et al., 2023; Hayes et al., 2025; Wang et al., 2025b; Madani et al., 2023; Xiao et al., 2024b; Su et al., 2024; Buehler and Buehler, 2024; Wang et al., 2025a; Zhu et al., 2024; Ferruz et al., 2022) have been highly successful because proteins/peptides are tokenized at the level of amino acid building blocks (which are akin to the words in a sentence). This makes it possible to associate the sequences of these building blocks with the fold and function of proteins, and then the corresponding building blocks can be assembled into targeted sequences.
Inspired from the successes of protein LLMs, LLMs have been increasingly applied in computational chemistry (Zhang et al., 2024a; b; Fang et al., 2023; Livne et al., 2024; Pei et al., 2023; Zhao et al., 2024; 2023; Yu et al., 2024; Ye et al., 2025; Liu et al., 2023b; Li et al., 2024c; Liu et al., 2024a; Taylor et al., 2022). Recent work has further explored cross-modal modeling of molecule and language (Edwards et al., 2022b; Liu et al., 2022; Li et al., 2025b; Chang and Ye, 2024). General-purpose LLMs such as GPT-4o (OpenAI, 2025a), GPT-5 (OpenAI, 2025b), Gemini-2.5-Flash (Google DeepMind, 2025), and Claude-3.5-H (Anthropic, 2025) show multimodal reasoning capabilities. However, most molecular language models still rely on atom-level (Wang et al., 2019; Ahmad et al., 2022; Zhou et al., 2023; Xia et al., 2023) or Byte-Pair Encoding on SMILES (Chang and Ye, 2024), and thus they often produce invalid tokens, misalign with chemical structures, and generate unsynthesizable molecules. To address these limitations, other approaches incorporate functional groups and fragments into molecular representations (Li et al., 2023a; Nguyen et al., 2024c; Han et al., 2023; Jin et al., 2024; Nguyen et al., 2024a; Wang et al., 2023; Zhang et al., 2020; 2021), capturing both atomic- and group-level information. However, these methods tend to break molecules at bonds that are difficult or even impossible to form through chemical reactions. In contrast, our mCLM decomposes small molecules into function-infused and synthesis-friendly building blocks and integrates synthesis constraints early in training, akin to the way peptide/protein LLMs tokenize at the amino acid level.
5 Conclusions and Future Work
We have developed a function- and synthesis-aware modular Chemical-Language Model (mCLM), as the first attempt to jointly model natural language sequences with modular chemical language. By design, the mCLM only generates chemical building blocks that can be iteratively assembled on robotic small molecule synthesis platforms, enabling the rapid creation of novel molecules with desired functions, all accessible by non-specialists. In the future we aim to incorporate richer multimodal knowledge related to physical and chemical properties from 2D/3D molecular structures, protein, cell lines and nucleic acid sequences to further enable the mCLM to reason on biological activity, protein docking knowledge, and individuals’ genetic profiles. We plan to extend chemical reasoning to additional aspects such as filling in unspoken knowledge gaps, thinking outside of the box, System 2 thinking for counterfactual reasoning and plausibility prediction, and resolving conflicting claims. We also plan to leverage more physical constraints from simulation tools and chemical and reaction knowledge bases. In the long term, we envision the mCLM as part of a comprehensive, multi-agent, human-in-the-loop autonomous laboratory, structured around iterative cycles of reasoning, proposal, synthesis, physical testing, feedback, and reasoning to enable never-ending self-improvement and co-evolvement with human scientists.
Reproducibility Statement
In section 3.1 and D.2.3, we describe the process and resources of developing oracle models for evaluation. The synthesis-guaranteed tokenizer is described in further detail in Section F. Sections C and  D explain the data collection process and the statistics of the training dataset. Section G lists all source datasets. Model training details are described in Section E. Evaluation data from FDA-approved drugs are described in Section 3.2, which also lists the baselines we used for comparison. The full data and resources will be released upon publication.
Appendix A Additional Results
A.1 Synthesizability
For the mCLM, all starting materials are the core building blocks, and the same collection of building blocks is used redundantly to make all the proposed structures. For this analysis, we assume that all the building blocks required for the mCLM approach are available. Given a sequence of blocks A - B - C - D in our methodology, we can assemble them with the following steps:
  • Coupling reaction (A and B)
  • Deprotect reaction (AB)
  • Coupling reaction (AB and C)
  • Deprotect reaction (ABC)
  • Coupling reaction (ABC and D)
Scaling this formula to $n$ building blocks gives a requirement of 2n32n-3 reactions (synthetic steps).
On the other hand, each molecule generated by MoleculeSTM is a unique problem for synthesis, and thus cannot leverage the strengths of the block-based approach. Each generated target requires a customized synthesis solution.
However, it is important to consider an external metric for verifying the synthesis capabilities of the mCLM. In table 3, we compute validity and synthetic accessibility scores for all baselines, both overall and in terms of specific property modifications. We find that mCLM outperforms all baselines (lower SA is better).
Table 3: Synthetic Accessibility (SA) scores with RDKit validity percentages across datasets.
Dataset FDA mCLM MoleculeSTM FineMolTex GPT-4o GPT-5 Gemini-2.5-F LDMol Claude-3.5-H
SA Valid (%) SA Valid (%) SA Valid (%) SA Valid (%) SA Valid (%) SA Valid (%) SA Valid (%) SA Valid (%) SA Valid (%)
AMES 2.70 100.0 2.39 100.0 2.64 94.26 2.52 94.26 3.06 90.98 2.98 72.13 2.92 89.34 2.85 90.16 3.12 93.44
BBBP 2.70 100.0 2.44 100.0 2.67 93.44 2.61 95.90 2.67 88.52 2.87 78.69 2.69 90.16 2.89 94.26 2.90 93.44
CYP3A4 2.70 100.0 2.40 100.0 2.60 95.90 2.63 94.26 2.70 89.34 2.83 85.25 2.75 93.44 2.74 93.44 2.69 88.52
DILI 2.70 100.0 2.38 100.0 2.62 94.26 2.59 93.44 2.73 93.44 2.75 81.15 2.70 90.16 2.80 87.70 2.79 88.52
HIA 2.70 100.0 2.42 100.0 2.73 91.80 2.59 93.44 2.68 87.70 2.82 73.77 2.77 94.26 2.80 90.16 2.86 91.80
PGP 2.70 100.0 2.45 100.0 2.63 93.44 2.59 94.26 2.81 90.16 2.85 86.07 2.79 82.79 2.87 93.44 2.75 90.16
Mean 2.70 100.0 2.41 100.0 2.64 93.85 2.59 94.26 2.78 90.02 2.85 79.51 2.77 90.03 2.83 91.53 2.85 90.98
A.1.1 More information on Retrosynthesis Experiments
Next, we select the best 3 models and consider a computationally expensive but rigorous approach. We leverage Allchemy, which is the state-of-the-art retrosynthesis software, to quantitatively measure this. We randomly sampled a representative set of 200 molecules generated by each baseline (and the original FDA molecules). We gave each molecule 30 minutes on a supercomputing cluster (which is quite exhaustive) and didn’t apply a price limit for substrates (so this is really quite generous for starting materials). Scores, as reported in Table 2, show much stronger overall performance of the mCLM. Here, validity is the percent of syntactically correct molecules as measured by RDKit, synthesizability is the percent of valid molecules where Allchemy could find a synthetic path, and Makeability is the product of those two metrics. Makeability represents the percent of generated molecules which could be made in the lab.
A.2 Expanded Testing Without Synthesis Guarantees
As a large-scale test, we apply the mCLM to improve all FDA-approved drugs consisting of at least 3 blocks (without synthesis guarantees). Note: this is just a confirmatory test, since the mCLM is intended to be used on the distribution of synthesis-guaranteed molecules (as shown in the main paper). This amounts to 430 molecular structures and 796 unseen blocks. Since most of these molecules (426/430) contain blocks that were not present in the 1,000 blocks used for training, this presents an opportunity to find how the mCLM performs out-of-distribution. Results in Table A.2 show that improvement is achieved for 5/6 properties, even though the model has not seen almost half of the blocks in its vocabulary during training. Specifically, the strictest synthesis-guaranteed tokenizer covers 26.7% of compounds in our corpus. For the remainder, a rule-based tokenizer is used. Despite this fallback, mCLM demonstrates strong generalization, as it successfully incorporated GNN-derived features for the unseen modules.
Table 4: Average pharmacokinetic and toxicity properties of FDA drugs with 3 or more blocks and their proposed modifications. Note, these molecules do not have synthesis-guarantees. ($\downarrow$: lower is better, $\uparrow$: higher is better.
Property AMES Mut. $(\downarrow)$ BBBP $(\uparrow)$ CYP3A4 Inhib. $(\downarrow)$ DILI $(\downarrow)$ HIA $(\uparrow)$ PGP $(\downarrow)$ Mean Improv. FDA Drug 59.5 37.6 2.0 66.2 93.2 66.0 0.00 % mCLM 54.0 41.4 1.2 55.5 97.6 68.0 12.87 %
A.3 Applying MoleculeSTM to the “Fallen Angels”
As a baseline comparison, we repeat the fallen angels reasoning experiment using MoleculeSTM (Figure 6). Even though the steps 1 and 2 of MoleculeSTM’s modification of TNG348 showed comparable property changes to mCLM, MoleculeSTM encounters molecule validity problems: it generates a syntactically incorrect molecule on step 1 of Evobrutinib and on step 3 of TNG348.
Figure 6: Examples of fallen angel property modification using MoleculeSTM.
Appendix B Critical Chemical Reasoning Algorithm
@@
8:  elseSelect property which requires most improvement.
9:  end if
@@
10:  mCLM generates a candidate modification $M^{\prime}$ by replacing, adding, or removing building blocks in $M$
11:  Evaluate $M^{\prime}$ with respect to desired functions
12:end while
Appendix C Pretraining Data Mixture